A sample of  Aluminium requires  4.0 J of energy to change its temp from 25 C to 32 C. What is the mass of the aluminum? Specific heat of Al = 0.89.




1
Expert's answer
2021-03-05T03:33:22-0500

"ΔE = mcΔT \\

m = \frac{ΔE}{cΔT} \\

ΔT = 32 -25 = 7 \;ºC \\

m = \frac{4.0}{0.89 \times 7} = 0.64 \;g"

Answer: 0.64 g